please answer 11 quickly
Answer #1

MENNINNE N RIWWW VODUN IAN.BRITISHUHADAPI P objective is to Po minimize :$ 6x + $5y constraints 11+ 24 €35 2n +3432 T PlottinSupplement 2 1 70+20 35 LLL + ub (Q10.7) +22+ 3y=32 (5,03 8 60,02 4 10 12 14 16 18 20 · Supplement iWWWwwse Finding corner point B : 70+ 2y = 35 - 0 aut 34 = 32- 3/ 7 n+24 = 35 21 an+34 = 32 __210 + 6/4 = 105 40+ y = 64 17 ~Find minobje minimum value of our constraint I At(0, 10:7) 06f0,32 Puthing this point in Gu+54- => 6.CO2+5(32) = 190 = 53-33DVA (160) Puting this point in but5y 64]+5[154] = 246+ = 1016 2121 PIPO 17 = 59. 76 so the point where costis minmur 3 is C5,